# Point3R: Streaming 3D Reconstruction with Explicit Spatial Pointer Memory

### v1

@@ Abstract
Dense 3D scene reconstruction from an ordered sequence or unordered image collections is a critical step when bringing research in computer vision into practical scenarios. Following the paradigm introduced by DUSt3R, which unifies an image pair densely into a shared coordinate system, subsequent methods maintain an implicit memory to achieve dense 3D reconstruction from more images. However, such implicit memory is limited in capacity and may suffer from information loss of earlier frames. We propose Point3R, an online framework targeting dense streaming 3D reconstruction. To be specific, we maintain an explicit spatial pointer memory directly associated with the 3D structure of the current scene. Each pointer in this memory is assigned a specific 3D position and aggregates scene information nearby in the global coordinate system into a changing spatial feature. Information extracted from the latest frame interacts explicitly with this pointer memory, enabling dense integration of the current observation into the global coordinate system. We design a 3D hierarchical position embedding to promote this interaction and design a simple yet effective fusion mechanism to ensure that our pointer memory is uniform and efficient. Our method achieves competitive or state-of-the-art performance on various tasks with low training costs. Code is available at: https://github.com/YkiWu/Point3R.
@@ 1 Introduction
Dense 3D reconstruction from image collections has long been a fundamental task in computer vision, with broad applications in fields such as autonomous driving [25, 77], medical modeling, and cultural heritage preservation. Conventional approaches [55, 1, 11, 67, 47, 66] first conduct a sequence of sub-tasks, including feature extraction and matching [33, 40], triangulation, and global alignment to get sparse geometry and camera poses. Multi-view stereo [20, 48, 22] is then used to obtain dense geometry. However, this tightly coupled pipeline is inefficient and tends to be vulnerable to noise. To address these challenges, DUSt3R [63] proposes a novel data-driven paradigm that directly aligns the input image pair and reconstructs it as point maps within a unified coordinate system.
Due to the constraint of pair-wise input, DUSt3R requires an additional global alignment step when performing dense reconstruction from multiple images, which is inefficient in multi-view settings. To improve this, subsequent works [70, 59, 58, 62] can be categorized into two main paradigms. One category of these works [70, 59] takes all input images simultaneously and employs global attention to reconstruct them into a unified coordinate system, requiring substantial computational resources and is misaligned with the incremental nature of real-world reconstruction scenarios. The second category [58, 62] introduces an external memory mechanism to retain information from past frames, enabling each new input to be directly integrated into the global coordinate system. For instance, Spann3R [58] maintains a memory that essentially caches implicit features of processed frames. However, this implicit memory often contains redundant information, and a discard strategy must be employed once memory capacity is saturated. CUT3R [62] uses a fixed-length token-based memory mechanism and directly updates the memory through interactions with image features. Nonetheless, as the number of processed frames increases, this fixed-capacity memory inevitably leads to the loss of earlier information.
@@ 2 Related Work
Conventional 3D Reconstruction: Classic approaches to 3D reconstruction from image collections are typically optimization-based and tailored to specific scenes. Structure-from-motion (SfM) [1, 11, 67, 66, 47, 55] follows a pipeline consisting of feature extraction [34, 45, 16], image matching [67, 31, 49, 9], triangulation to 3D, and bundle adjustment [2, 57] to obtain sparse geometry and estimated camera poses. Building upon this, subsequent methods such as Multi-view Stereo (MVS) [19, 48, 18, 64], Neural Radiance Fields (NeRF) [36, 8, 37, 61, 4], and 3D Gaussian Splatting (3DGS) [26] leverage known camera parameters to recover dense geometry or high-fidelity scene representation. These approaches rely on a sequential combination of multiple modules, which not only require considerable optimization time but are also vulnerable to noise. It is worth noting that Simultaneous Localization and Mapping (SLAM) [13, 17, 39, 27] can perform localization and reconstruction in an online manner. However, they often rely on specific camera motion assumptions [39, 13] (sometimes these motion assumptions may be misleading or restrictive) or require additional depth/LiDAR sensors [38] for better performance.
@@
Learning-Based 3D Reconstruction: To enhance accuracy and efficiency, the field of 3D reconstruction is gradually shifting toward learning-based and end-to-end paradigms. Some approaches utilize learnable modules to replace hand-crafted components [14, 46] during the traditional reconstruction pipeline. Some others try to optimize the overall pipeline in an end-to-end manner [56, 60, 71]. DUSt3R [63] introduces a pointmap representation and directly learns to integrate an image pair into the same coordinate system, which unifies all sub-tasks we have mentioned above. MonST3R [74] extends this paradigm to dynamic scenes by fine-tuning it on dynamic datasets. The pair-wise formulation needs a global alignment to process more views. Subsequent works [58, 62, 70, 59] are exploring how to further replace the global optimization step with an end-to-end learning framework. These efforts can be broadly categorized into two main branches. The former [70, 59] feeds all images simultaneously and leverages global attention to reconstruct the scene within a unified coordinate system; the latter [58, 62] proposes a streaming paradigm, in which a memory module stores information from past frames, thereby enabling online incremental reconstruction from sequential inputs. The streaming paradigm aligns more closely with real-world applications, offering improved scalability without imposing excessive computational demands.
@@
Streaming Reconstruction and Memory Mechanism: The streaming reconstruction paradigm and the memory mechanism are inherently aligned with each other. Existing streaming reconstruction methods [17, 10, 75, 58, 62] universally incorporate a certain form of memory to store information from past frames. This memory can take on various forms, such as explicit scene representation (the most direct form of memory), recurrent neural network architectures [10], and encoded or learnable token features [58, 62]. Explicit 3D scene representation is compact and efficient but tailored to a specific scene, limiting its generalizability. Spann3R [58] stores implicit features from previous frames in the memory, which may lead to redundancy. CUT3R [62] employs a fixed-length set of learnable token features as its memory module, which is continuously updated during sequential processing. However, its limited capacity can lead to information loss. In contrast, we propose a spatial pointer memory, in which each pointer is dynamically assigned a 3D position. This ensures that the total amount of stored information scales naturally with the extent of the explored scene. Furthermore, each pointer has a spatial feature that captures aggregated scene information nearby.
@@ 3.2 Pointer-Image Interaction
Interaction decoders. Our spatial pointer memory $\mathcal{M}$ consists of a set of 3D pointers (3D positions ${P}$ and spatial features ${M}$). Before processing the first frame, the memory has not yet stored any global spatial features of the current scene. Therefore, we use a simple layer to embed the image tokens ${F}_{0}$ of the first frame, and use the output ${M}_{0}$ to initialize the features of our memory. It is worth noting that, since the precise spatial positions represented by these features are not yet known at this time, each feature has not been assigned a specific 3D position. Then we use two intertwined decoders [65, 63] to enable interaction between the current image tokens and the memory:
@@
where we use a learnable pose token ${z}$ as a bridge between the current frame and the global coordinate system. After this interaction, we get the updated image tokens ${F^{\prime}_{t}}$ and pose token ${z^{\prime}_{t}}$. Then we use ${F^{\prime}_{t}}$ and ${z^{\prime}_{t}}$ to predict two pointmaps (${\hat{X}_{t}}^{self}$ in the local coordinate system of the current input and ${\hat{X}_{t}}^{global}$ in the global coordinate system) with their own confidence maps (${C_{t}}^{self}$and ${C_{t}}^{global}$), and a camera pose ${\hat{T}_{t}}$ representing the rigid transformation between this two coordinate system:
@@
where ${R}_{u,v}$ is the corresponding image patch and we use ${\hat{X}}_{t-1}^{global}$ from $t-1$ because we assume that the image tokens of the current frame are more likely to be spatially close to those of the previous frame. Of course, even in cases of significant pose shifts or unordered inputs, this assumption will not introduce any adverse effects. This is because the image tokens of the current frame interact with all memory features in our interaction decoders, and our hierarchical position embedding merely provides a potential prior. Due to the space limit, we will add more details about the design and implementation of our 3D hierarchical position embedding in the supplementary material.
@@ 3.3 Training Strategy
Training stages. Our model is trained in three stages. We train the model by sampling 5 frames per sequence in the first stage. The input here is 224$\times$224 resolution. Then we use input with different aspect ratios (set the maximum side to 512) in the second stage, following CUT3R [62]. And finally, we freeze the encoder and fine-tune other parts on 8-frame sequences.
Implementation details. We initialize our ViT-Large [63, 15] image encoder, ViT-Base interaction decoders [63, 65], and DPT [42] heads with pre-trained weights from DUSt3R [63]. Our memory encoder is composed of a light-weight ViT (6 blocks) and a 2-layer MLP. Each memory feature has a dimensionality of 768. We use the AdamW optimizer [32] and the learning rate warms up to a maximum value of 5e-5 and decreases according to a cosine schedule. We train our model on 8 H100 NVIDIA GPUs for 7 days, which is a relatively low cost.
@@ 4.2 Monocular and Video Depth Estimation
Method Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$
@@
Ours 0.079 92.0 0.395 56.8 0.061 95.4 0.087 93.7
Video Depth Estimation. We align predicted depth maps to ground truth using a per-sequence scale (Per-sequence alignment) to evaluate per-frame quality and inter-frame consistency. We also compare results without alignment with other metric pointmap methods like MASt3R [29] and CUT3R [62] (Metric-scale alignment). As shown in Table 3, with the per-sequence scale alignment, our method outperforms DUSt3R [63], MASt3R [29], and Spann3R [58] by a large margin. These methods have a static scene assumption and are trained only on static datasets. Our spatial pointer memory directly associates spatial features with their real-world positions and imposes no assumptions or dependencies on whether the scene is static or dynamic. Our method performs comparably, or even better than MonST3R-GA [74] and CUT3R [62] (methods trained on dynamic datasets). Besides, in the metric-scale setting, our method outperforms MASt3R-GA [29] and performs comparably with CUT3R [62], leading on Sintel and Bonn.
@@
Alignment Method Optim. Onl. Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$<$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$ <$1.25\uparrow$
@@ 4.4 Analysis and Discussion
Effect of the 3D hierarchical position embedding. We report the 3D reconstruction results without the 3D hierarchical position embedding in Table 5, which shows the effectiveness of the elaborate position embedding we proposed.

### v2

@@ Abstract
Dense 3D scene reconstruction from an ordered sequence or unordered image collections is a critical step when bringing research in computer vision into practical scenarios. Following the paradigm introduced by DUSt3R, which unifies an image pair densely into a shared coordinate system, subsequent methods maintain an implicit memory to achieve dense 3D reconstruction from more images. However, such implicit memory is limited in capacity and may suffer from information loss of earlier frames. We propose Point3R, an online framework targeting dense streaming 3D reconstruction. To be specific, we maintain an explicit spatial pointer memory directly associated with the 3D structure of the current scene. Each pointer in this memory is assigned a specific 3D position and aggregates scene information nearby in the global coordinate system into a changing spatial feature. Information extracted from the latest frame interacts explicitly with this pointer memory, enabling dense integration of the current observation into the global coordinate system. We design a 3D hierarchical position embedding to promote this interaction and design a simple yet effective fusion mechanism to ensure that our pointer memory is uniform and efficient. Our method achieves competitive or state-of-the-art performance on various tasks with low training costs. Code: https://github.com/YkiWu/Point3R.
@@ 1 Introduction
Dense 3D reconstruction from image collections has long been a fundamental task in computer vision, with broad applications in fields such as autonomous driving [25, 78], medical modeling, and cultural heritage preservation. Conventional approaches [56, 1, 11, 68, 48, 67] first conduct a sequence of sub-tasks, including feature extraction and matching [33, 41], triangulation, and global alignment to get sparse geometry and camera poses. Multi-view stereo [20, 49, 22] is then used to obtain dense geometry. However, this tightly coupled pipeline is inefficient and tends to be vulnerable to noise. To address these challenges, DUSt3R [64] proposes a data-driven paradigm that directly reconstructs the input image pair as point maps within a unified coordinate system.
Due to the constraint of pair-wise input, DUSt3R requires an additional global alignment step when performing dense reconstruction from multiple images, which is inefficient in multi-view settings. To improve this, subsequent works [71, 60, 59, 63] can be categorized into two main paradigms. One category of these works [71, 60] takes all input images simultaneously and employs global attention to reconstruct them into a unified coordinate system, requiring substantial computational resources and is misaligned with the incremental nature of real-world reconstruction scenarios. The second category [59, 63] introduces an external memory mechanism to retain information from past frames, enabling each new input to be directly integrated into the global coordinate system. For instance, Spann3R [59] maintains a memory that essentially caches implicit features of processed frames. However, this implicit memory often contains redundant information, and a discard strategy must be employed once capacity is saturated. CUT3R [63] uses a fixed-length token-based memory mechanism and directly updates it through interactions with image features. Nonetheless, as the number of processed frames increases, this memory inevitably leads to the loss of earlier information.
@@ 2 Related Work
Conventional 3D Reconstruction. Classic approaches to 3D reconstruction from image collections are typically optimization-based and tailored to specific scenes. Structure-from-motion (SfM) [1, 11, 68, 67, 48, 56] follows a pipeline consisting of feature extraction [34, 46, 16], image matching [68, 31, 50, 9], triangulation to 3D, and bundle adjustment [2, 58] to obtain sparse geometry and estimated camera poses. Building upon this, subsequent methods such as Multi-view Stereo (MVS) [19, 49, 18, 65], Neural Radiance Fields (NeRF) [36, 8, 37, 62, 4], and 3D Gaussian Splatting (3DGS) [26] leverage known camera parameters to recover dense geometry or high-fidelity scene representation. These approaches rely on a sequential combination of multiple modules, which not only require considerable optimization time but are also vulnerable to noise. It is worth noting that Simultaneous Localization and Mapping (SLAM) [13, 17, 40, 27] can perform localization and reconstruction in an online manner. However, they often rely on specific camera motion assumptions [40, 13] (sometimes these motion assumptions may be misleading or restrictive) or require additional depth/LiDAR sensors [39] for better performance.
@@
Learning-Based 3D Reconstruction. To enhance accuracy and efficiency, the field of 3D reconstruction is gradually shifting toward learning-based and end-to-end paradigms. Some approaches utilize learnable modules to replace hand-crafted components [14, 47] during the traditional reconstruction pipeline. Some others try to optimize the overall pipeline in an end-to-end manner [57, 61, 72]. DUSt3R [64] introduces a pointmap representation and directly learns to integrate an image pair into the same coordinate system, which unifies all sub-tasks we have mentioned above. MonST3R [75] extends this paradigm to dynamic scenes by fine-tuning it on dynamic datasets. However, this pair-wise formulation needs a global alignment to process more views, which is computationally intensive and time-consuming. Subsequent works [59, 63, 71, 60] are exploring how to further replace the global optimization step with an end-to-end learning framework. These efforts can be broadly categorized into two main branches. The former [71, 60] feeds all images simultaneously and leverages global attention to reconstruct the scene within a unified coordinate system; the latter [59, 63] proposes a streaming paradigm, in which a memory module stores information from past frames, thereby enabling online incremental reconstruction from sequential inputs. The streaming paradigm aligns more closely with real-world applications, offering improved scalability without imposing excessive computational demands.
@@
Streaming Reconstruction and Memory Mechanism. The streaming reconstruction paradigm and the memory mechanism are inherently aligned with each other. Existing streaming reconstruction methods [17, 10, 76, 59, 63] universally incorporate a certain form of memory to store information from past frames. This memory can take on various forms, such as explicit scene representation (the most direct form of memory), recurrent neural network architectures [10], and encoded or learnable token features [59, 63]. Explicit 3D scene representation is compact and efficient but tailored to a specific scene, limiting its generalizability. Spann3R [59] stores implicit features from previous frames in the memory, which may lead to redundancy. CUT3R [63] employs a fixed-length set of learnable token features as its memory module, which is continuously updated during sequential processing. However, its limited capacity can lead to information loss. In contrast, we propose a spatial pointer memory, in which each pointer is dynamically assigned a 3D position. This ensures that the total amount of stored information scales naturally with the extent of the explored scene. Furthermore, each pointer has a spatial feature that captures aggregated scene information nearby.
@@ 3.2 Pointer-Image Interaction
Interaction Decoders. Our spatial pointer memory $\mathcal{M}$ consists of a set of 3D pointers (3D positions P{P} and spatial features M{M}). Before processing the first frame, the memory has not yet stored any global spatial features of the current scene. Therefore, we use a simple layer to embed the image tokens ${F}_{0}$ of the first frame, and use the output ${M}_{0}$ to initialize the features of our memory. It is worth noting that, since the precise spatial positions represented by these features are not yet known at this time, each feature has not been assigned a specific 3D position. Then we use two intertwined decoders [66, 64] to enable interaction between the current image tokens and the memory:
@@
where we use a learnable pose token z{z} as a bridge between the current frame and the global coordinate system. After this interaction, we get the updated image tokens $F^{\prime}_{t}$} and pose token $z^{\prime}_{t}$}. Then we use $F^{\prime}_{t}$} and $z^{\prime}_{t}$} to predict two pointmaps (${\hat{X}_{t}}^{self}$ in the local coordinate system of the current input and ${\hat{X}_{t}}^{global}$ in the global coordinate system) with their own confidence maps (${C_{t}}^{self}$and ${C_{t}}^{global}$), and a camera pose $\hat{T}_{t}$} representing the rigid transformation between this two coordinate system:
@@
where ${R}_{u,v}$ is the corresponding image patch and we use ${\hat{X}}_{t-1}^{global}$ from t1t-1 because we assume that the image tokens of the current frame are more likely to be spatially close to those of the previous frame. Of course, even in cases of significant pose shifts or unordered inputs, this assumption will not introduce any adverse effects. This is because the image tokens of the current frame interact with all memory features in our interaction decoders, and our hierarchical position embedding merely provides a potential prior. Due to the space limit, we will add more details about the design and implementation of our 3D hierarchical position embedding in the supplementary material.
@@ 3.3 Training Strategy
Training Stages. Our model is trained in three stages. We train the model by sampling 5 frames per sequence in the first stage. The input here is $\times224$ resolution. Then we use input with different aspect ratios (set the maximum side to 512) in the second stage, following CUT3R [63]. And finally, we freeze the encoder and fine-tune other parts on 8-frame sequences.
Implementation Details. We initialize our ViT-Large [64, 15] image encoder, ViT-Base interaction decoders [64, 66], and DPT [43] heads with pre-trained weights from DUSt3R [64]. Our memory encoder is composed of a light-weight ViT (6 blocks) and a 2-layer MLP. Each memory feature has a dimensionality of 768. We use the AdamW optimizer [32] and the learning rate warms up to a maximum value of 5e-5 and decreases according to a cosine schedule. We train our model on 8 A800 NVIDIA GPUs for 15 days, which is a relatively low cost.
@@ 4.2 Monocular and Video Depth Estimation
Method Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$$1.25\uparrow$
@@
Ours 0.078 92.3 0.395 55.0 0.061 94.5 0.083 94.6
Video Depth Estimation. We align predicted depth maps to ground truth using a per-sequence scale (Per-sequence alignment) to evaluate per-frame quality and inter-frame consistency. We also compare results without alignment with other metric pointmap methods like MASt3R [29] and CUT3R [63] (Metric-scale alignment). As shown in Table 3, with the per-sequence scale alignment, our method outperforms DUSt3R [64], MASt3R [29], and Spann3R [59] by a large margin. These methods have a static scene assumption and are trained only on static datasets. Our spatial pointer memory directly associates spatial features with their real-world positions and imposes no assumptions or dependencies on whether the scene is static or dynamic. Our method performs comparably, or even better than MonST3R-GA [75] and CUT3R [63] (methods trained on dynamic datasets). Besides, in the metric-scale setting, our method outperforms MASt3R-GA [29] and performs comparably with CUT3R [63], leading on Bonn.
@@
Alignment Method Optim. Onl. Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$$1.25\uparrow$ Abs Rel $\downarrow$ $\delta$ $1.25\uparrow$
@@ 4.4 Analysis and Discussion
Comparison on Long Sequences. We compare our model with CUT3R [63] on long sequences to show the advantage of our explicit spatial pointer memory to effectively store more information from past frames. We resampled the testing sequences from 7-scenes at the interval of 1 (each contains 500-1000 frames) and NRGBD at the interval of 2 (400-900 frames). As shown in the Table 5, our method outperforms CUT3R by a large margin, showing the advantage of our model when handling long sequences in practical applications.
Robustness to the Input Orders. We evaluate the robustness of our model to the order of camera registration. To be specific, for each scene in 7-scenes, we sample a sequence with a frame interval of 20 (each sequence contains 25-50 frames). For each scene in NRGBD, we sample a sequence with a frame interval of 40 (each sequence contains 20-40 frames). Then we disrupt the input sequence after the sampling to evaluate the reconstruction results. From Table 6, we can observe that our method still achieves good reconstruction performance after the input sequence is shuffled. This verifies the robustness of our explicit pointer memory to handle the discontinuity of the inputs.
Effect of the Memory Fusion Mechanism and Runtime Analysis. We design the memory fusion mechanism to get a balance between efficiency and performance. Figure 4 shows the changes in the number of pointers $\mathrm{N}$ in the memory (line graph) and per-frame runtime $\mathrm{t}$ (histogram) with the increasing number of frames $\mathrm{K}$ when processing Scene $\mathit{WhiteRoom}$ (from NRGBD dataset). We see that this memory fusion mechanism can control the total number of pointers and per-frame runtime within a reasonable range. We also report the 3D reconstruction results without the memory fusion mechanism on the 7-scenes and NRGBD datasets in Table 7. Although this fusion mechanism may result in a slight decrease in some metrics, we believe that efficiency improvement is worthwhile.
Effect of the 3D Hierarchical Position Embedding. To demonstrate the effectiveness of our 3D hierarchical position embedding, we conduct ablation studies for this module in Table 7. We report the 3D reconstruction results without the 3D hierarchical position embedding (“w/o 3DHPE”), which shows the effectiveness of the elaborate position embedding we proposed.
Table 5: Quantitative 3D reconstruction results on long sequences.
7-scenes NRGBD
@@
Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med.
CUT3R [63] 0.238 0.172 0.105 0.025 0.527 0.537 0.372 0.270 0.211 0.090 0.556 0.582
Ours 0.071 0.033 0.031 0.015 0.558 0.587 0.110 0.050 0.025 0.009 0.641 0.729
Table 6: Robustness to the input orders.
7-scenes NRGBD
[…]
Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med.
Ordered 0.032 0.016 0.024 0.008 0.665 0.758 0.064 0.031 0.029 0.011 0.801 0.949
Shuffled 0.033 0.014 0.019 0.010 0.669 0.764 0.063 0.027 0.029 0.009 0.800 0.953
@@
Comparison with SLAM Systems and More Discussions. Traditional SLAM systems and the recent feed-forward 3D reconstruction methods (e.g., DUSt3R [64], CUT3R [63], and ours) are similar in terms of output format (reconstruction results and predicted poses), but their core objectives are different. The main objective of SLAM lies in accurately registering the camera poses from RGB-D inputs (where the depth could be obtained from prediction models). Differently, feed-forward 3D reconstruction methods aim to develop a unified model to predict 3D reconstructed points for each frame in a shared coordinate system (and some recent methods simultaneously predict the camera pose with an additional head). To better evaluate these two paradigms, we compare our method with MASt3R-SLAM [38] on its own established benchmarks.
We conduct the dense geometry evaluation on 7-Scenes seq-01 and also report the RMSE of the absolute trajectory error (ATE) in meters. We also compare these two methods in terms of peak memory usage and per-frame runtime. As shown in Table 8, our method achieves better dense geometry reconstruction results with lower memory consumption. However, there is still room for further improvement in tracking accuracy (as discussed in Limitations) and runtime efficiency. We hold the belief that these two paradigms are compatible and essentially beneficial to each other. Visual SLAM systems could use a feed-forward 3D reconstruction model to produce reconstructed points (just like MASt3R-SLAM [38]). Feed-forward 3D reconstruction methods can further employ SLAM techniques such as bundle adjustment to obtain more accurate poses.
Table 8: Comparison with SLAM in mapping quality, tracking accuracy and efficiency.
[…]
MASt3R-SLAM [38] 0.068 0.045 0.066 7.18 GB 0.11 s
Ours 0.061 0.022 0.084 5.46 GB 0.20 s